At a school supply store, a saleswoman sold 32 notebookes on Monday, some for $1.50 each and the rest for $2.00 each. If her notebook receipts that day totaled $51.00, how many of each type of notebook did she sell?
Answer by stanbon(75887)

At a school supply store, a saleswoman sold 32 notebookes on Monday, some for $1.50 each and the rest for $2.00 each. If her notebook receipts that day totaled $51.00, how many of each type of notebook did she sell?
---------
Quantity Eq:::: a + b = 32
Value equation:1.5a+2b = 51 dollars
------------------------------------
Multiply thru the 1st equation by 20
Multiply thru the 2nd equation by 10
--------
20a + 20b = 20*32
15a + 20b = 10*51
------------------------------
Subtract and solve for "a":
5a = 130
a = 26 (# of $1.50 notebooks sold)
---
Since a+b=32, b = 6 (# of $2 notebooks sold)
------------------
Cheers,
Stan H.
------------------